An object is thrown vertically up from the ground passes the height 5 m twice in an interval of 10 secs. what is the time of flight?

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\T_{flight}={2 v \over g} \\=>v={gT_{flight} \over 2} \\=>v=50m/s $ at height 5m$ \\v^2=50^{2}+2gs=2500+100=2600 \\=>v=\sqrt{2600} \\=>T={2 v \over g}=2\sqrt{26} \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty